OpenStudy (anonymous):

Please Help Me solving this problem! -2(5x (-10/7)-1)-(-10/7)=-4(-10/7 -3)

OpenStudy (blurbendy):

start with the inner most parentheses first on the left side, which is distributing the 5x to -10/7. what does that give you?

OpenStudy (anonymous):

it is supposed to be \[5\times -\frac{ 10 }{ 7}\]

OpenStudy (blurbendy):

so it's -2(5x-10/7)-1)-(-10/7)=-4(-10/7 -3) ???

OpenStudy (blurbendy):

or is the x supposed to be multiplication, not a variable

OpenStudy (anonymous):